[figure: 9.7 cm]

SEMICONDUCTING PROPERTIES OF DISORDERED HALF-HEUSLER SYSTEMS

J. Toboła^{a}, G. Venturini ^{b} and S. Kaprzyk^{a}

aFaculty of Physics and Nuclear Techniques, University of Mining and Metallurgy,

30-059 Cracow, Poland bLaboratoire de Chimie du Solide Mineral, Université H. Poincaré, Vandoeuvre-les-Nancy, 54042 Nancy Cedex, France

Recently, a composition induced metal-semiconductor-metal crossover has been detected in the Fe_{1−x}Ni_{x}TiSb half-Heusler system using resistivity and thermopower measurements and supported by electronic structure computations based on the KKR-CPA method [1]. Such an unusual behavior is also expected in other disordered half-Heusler phases possessing the critical number of the valence electrons. We present results of electronic structure calculations on Fe_{1−x}Pt_{x}ZrSb, Ru_{1−x}Ni_{x}ZrSb and Ru_{1−x}Pd_{x}ZrSb which are isoelectronic to the aforementioned system. The onset of semiconducting state, when alloying parent compounds (metals), is detected through a remarkable modification of density of states as a function of composition. Moreover, a shift of the Fermi level (E_{F}) from valence to conduction states for x ≈ 0.5, should have an apparent effect on conductivity and thermopower behavior. The appearance of the energy gap at E_{F} in the disordered alloys is also in line with previous studies on ordered half-Heusler systems, which revealed a high propensity for forming bandgaps in the compounds having 18 valence electrons.

[1] J. Toboła et al., Phys. Rev. B 64, 155103 (2001).

[figure: 13.4 cm]

Subject category :

7. Computations of Electronic Band Structures

Presentation mode : oral

Corresponding author : J. Toboła

Address for correspondence :

1. Faculty of Physics and Nuclear Techniques, AGH Al. Mickiewicza 30, 30-059 Cracow, Poland

2. Laboratoire de Physique des Materiaux, Ecole des Mines Parc du Saurupt, 54042 Nancy Cedex, France

Email address : tobola@ftj.agh.edu.pl